OpenStudy (anonymous):

now i need help to solve: 3d + 4 = 2 + 3d - 1/2 need to show my work too. #help

OpenStudy (saifoo.khan):

add like terms by changing their sides. when u are going to change that, the signs of terms will change as well.

OpenStudy (anonymous):

so 3d + 3d?

OpenStudy (anonymous):

? :)

OpenStudy (saifoo.khan):

NO.

OpenStudy (anonymous):

then whatt

OpenStudy (saifoo.khan):

3d + 4 = 2 + 3d - 1/2 3d - 3d = 2 - 1/2 - 4 0 = -2 -1/2 0 = -3/2

OpenStudy (anonymous):

thanks <3

OpenStudy (saifoo.khan):

np